# A $1.9\dot{9}$-approximation algorithm for the vertex cover problem on undirected, unweighted graphs(?)

### v1

@@ Abstract
The DiRe (Diverse + Representative) committee feasibility problem (a.k.a. the minimum vertex cover problem on unweighted undirected graphs) concerns the determination of the smallest size committee that satisfies the given constraints. Here, for this problem, we discover an unconditional deterministic polynomial-time algorithm that is an amalgamation of maximum matching, breadth-first search, maximal matching, and local minimization.
Preface: The DiRe committee feasibility problem (stated in the abstract) and the vertex cover problem on unweighted undirected graphs are equivalent (vertices $=$ candidates; edges $=$ candidate groups / voter populations’ approved candidates; for details, see Appendix A). Hence, for technical simplicity, we henceforth focus the discussion on the latter problem.
@@ 1 Introduction
Given an unweighted undirected graph (specifically, a 2-uniform hypergraph), the vertex cover of the graph is a set of vertices that includes at least one endpoint of every edge of the graph. Formally, given a graph $G=(V,E)$ consisting of a set of vertices $V$ and a collection $E$ of 2-element subsets of $V$ called edges, the vertex cover of the graph $G$ is a subset of vertices $S\subseteq V$ that includes at least one endpoint of every edge of the graph, i.e., for all $e\in E$, $e\cap S\neq\phi$. The corresponding computational problem of finding the minimum-size vertex cover (MVC) is NP-complete [^3] Strictly speaking, the decision version of the vertex cover problem is NP-complete whereas MVC itself (search version) is NP-hard. See Section 2.1 of [Kho19] for a lucid explanation delineating (a) search and decision problems and (b) NP-hardness and NP-completeness. [Coo71, Lev73, Kar72], which means that there is no known deterministic polynomial-time algorithm to solve MVC. Here, we present an unconditional deterministic polynomial-time algorithm for MVC on unweighted simple connected graphs [^4] We subtly yet drastically switch the discussion from unweighted undirected graphs to unweighted simple connected graphs. For simplicity, we want to avoid having loops and/or unconnected components in the graph. In the context of this paper, this switch has no impact on the NP-completeness of the problem (Appendix C). Notwithstanding, in the case of the presence of loops, our algorithm will work (with minor modifications) if each loop is replaced by adding a dummy vertex and a corresponding edge. In the case of unconnected components, we can run the algorithm for each connected component independently and take a union of each of the minimum vertex covers to get the final minimum vertex cover..
@@ 3.3 Maximal Matching
Given a graph $G$, a matching $M$ is said to be maximal if for all other matching $M^{\prime}$, $M\not\subset M^{\prime}$.
@@
We specifically use Lemma 3 in Section 5 and explain why the third phase is called maximal matching and not maximum matching.
@@ 3.5 Summary
The algorithm we discovered is an amalgamation of the above-discussed phases. The sequential implementation of these phases ensures we get a minimum vertex cover cover. At a high-level, this is because: (i) Maximum matching and breadth-first search ensures that the edges selecting during the maximal matching phase follows a procedure as opposed to vanilla maximal matching where edges are selected randomly. (ii) Maximal matching implies we get a vertex cover. (iii) Local minimization ensures we get the smallest vertex cover. Overall, the combination of all these implies we get the minimum vertex cover.
Non-technical Comment: As discussed, the flow of the algorithm is as follows: maximum matching normal-$\rightarrow$ breadth-first search normal-$\rightarrow$ maximal matching normal-$\rightarrow$ local minimization. However, the evolution of the algorithm happened in the following order: maximal matching normal-$\rightarrow$ local minimization normal-$\rightarrow$ breadth-first search normal-$\rightarrow$ maximum matching. Indeed, eventually “prefixing” the algorithm with maximum matching helped us deal with the messy cycles, especially odd cycles. Recall that Blossom algorithm [Edm65] had to do “extra work” just to deal with odd cycles.
@@ 9       $S$ = $\mathtt{Local\_Minimization}$($R$)
Result: returns $R$ Represents Table
1 $R$ = a two-column table, Represents Table, that stores the endpoints of an edge selected during maximal matching and the corresponding vertices each endpoint represents
2 for each $level$ in $BFS_{level}$ do
@@ 16             If any vertex becomes edgeless in $G$, mark the vertex as visited in $BFS_{level}$
Result: returns $S$ the smallest vertex cover
1 $S$ = $\phi$
@@ 6             $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $v$, $\phi$)
11// The following for loop will traverse through the table $R$ top-down
12 for each $row$ in $R$ do
13       if if any one endpoint in $row$ is either frozen or removed then
@@
15       else if one endpoint $u$ in $row$ only represents another endpoint vertex $v$ in $row$ and $v$ represents more than one vertex then
@@ 17                   $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $v$, $u$)
24// The following for loop will traverse through the table $R$ bottom-up
25 for each $row$ in $R$ do
@@ 29             $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $u$, $\phi$)
32             //at this point, both endpoints $u$ and $v$ in $row$ represent exactly one vertex, namely each other
@@ 40                   $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $u$, $v$)
Data: Represents Table $R$, Vertex Cover $S$, vertex to be frozen $freeze$, vertex to be removed $remove$
@@
1 Remove vertex $remove$ and its represents list from $R$
2 Freeze vertex $freeze$ in $R$
3 Append vertex $freeze$ to $S$
4 Remove vertex $freeze$ from every represents list in $R$
5 Remove the represents list of vertex $freeze$ in $R$
6 for each non-frozen and unremoved $endpoint$ in $R$ that represents $remove$ do
7       $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $endpoint$, $\phi$)
@@
9for each non-frozen and unremoved $endpoint$ in $R$ that does not represent any vertex do
10       $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $\phi$, $endpoint$)
@@
12return $R$, $S$
Algorithm 4 $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $freeze$, $remove$)
5 Proof of Correctness
In this section, we show the correctness of the algorithm by proving the following theorem:
Theorem 1.
@@ 5 Proof of Correctness
We prove the theorem through a sequence of lemmas. Foremost, in the forward direction, we have the following lemma:
Lemma 4.
If a given instance of VC is a “Yes” instance, then the Algorithm 1 returns “Yes”.
Proof.
If the given instance of VC is a “Yes” instance, then Line 4 of Algorithm 1 can never return “No” as $k\geq|E_{M}|$ (Lemma 1). Also, the execution will never reach Line 14 of Algorithm 1 as Line 11 will return “Yes” during one of the $m$ iterations when Algorithm 3 finds the minimum vertex cover because $k\geq|S|$. ∎
Next, in the reverse direction, we have the following lemma:
Lemma 5.
If the Algorithm 1 returns “Yes”, then the given instance of VC is a “Yes” instance.
We prove Lemma 5 in the following sequence: (i) we prove that variable $S$ in Line 9 of Algorithm 1 is always a vertex cover (Lemma 6), (ii) we prove that this vertex cover is the smallest vertex cover based on the endpoints selected during maximal matching (Lemma 7), and (iii) there exists at least one minimum vertex cover among the $m$ smallest vertex covers (Lemma 8).
Lemma 6.
Variable $S$ in Line 9 of Algorithm 1 is a vertex cover.
Proof.
Algorithm 2 is an algorithm for maximal matching. The endpoints of the edges selected during maximal matching form a vertex cover (Lemma 2). However, Algorithm 3 removes some of these endpoints. But an endpoint is removed only if every endpoint that represents the removed endpoint is (i) already frozen [^6] Recall that a frozen vertex is implicitly always added to the set of vertices $S$. or (ii) is immediately frozen. This is equivalent to ensuring that each edge has at least one endpoint in the set of vertices. This implies that the set of vertices returned by Algorithm 3 is a vertex cover. Hence, variable $S$ in Line 9 of Algorithm 1 is a vertex cover. ∎
Lemma 7.
Given a set of endpoints $V^{\prime}\subseteq V$ of edges selected during maximal matching, variable $S$ in Line 9 of Algorithm 1 is the smallest vertex cover such that for all vertex covers $S^{\prime}\subseteq V^{\prime}$, $|S|\leq|S^{\prime}|$.
Proof.
Based on Lemma 6, $S$ is guaranteed to be a vertex cover. Hence, it remains to be proven that $S$ is the smallest vertex cover derivable from the endpoints selected during maximal matching (Algorithm 2). To do so, we use a couple of observations:
Observation 1.
Given a represents table $R$ consisting of $r$ rows, an endpoint vertex $v$ in row $i$ cannot represent an endpoint vertex $u$ in row $j$, for all $j<i$ where $i,j\in\mathbb{N}$ and $1\leq i,j\leq r$.
When an edge is selected during maximal matching, all the edges covered by two of its endpoints are removed. Simultaneously, the represents table $R$ is updated to reflect the two endpoints and the vertices each of the endpoints represent (Line 14 - Algorithm 2). Hence, the succeeding entry in $R$ can not represent any of the endpoints already present in the table $R$.
Next observation is related to the distance between (i) an endpoint and the vertices it represents and (ii) the endpoint and the endpoints it is represented by. Such distance is at most one level in the $BFS_{level}$.
Observation 2.
Given a represents table $R$ and a BFS level table $BFS_{level}$, an endpoint vertex $v$ in $R$ at level $i$ in $BFS_{level}$ can represent or can be represented by a vertex $u$ that is at level $i-1$, $i$ or $i+1$ in $BFS_{level}$.
Finally, we proceed to show that variable $S$ in Line 9 of Algorithm 1 is the smallest vertex cover such that for all vertex covers $S^{\prime}\subseteq V^{\prime}$, $|S|\leq|S^{\prime}|$. To do so, we prove that Algorithm 3 returns the smallest vertex cover derivable from the endpoints selected during maximal matching (Algorithm 2).
  • Lines 2 to 8 of Algorithm 3 freezes “necessary” vertices: Variable $P$ (Line 2 - Algorithm 3) consists of the endpoints of the edges selected by Algorithm 2. Any endpoint $v$ in represents table $R$ that represents a vertex not in $P$ needs to be frozen and added to the vertex cover $S$. If a vertex $u$ is not in $P$, it implies that, by design, it will not be in the vertex cover. Hence, any vertex that is connected to $u$ via an edge will be an endpoint in $R$ and in turn, it needs to be in the vertex cover.
  • Lines 10 to 18 of Algorithm 3 removes “terminal” vertices: A terminal vertex $u$ in a connected graph does not bring value to the vertex cover because (i) it represents one vertex $v$ and (ii) it is represented by one vertex $v$. On the other hand, given that the graph $G$ is connected and consists of more than two vertices [^7] In case the graph $G$ has two vertices, it is a trivial case and any one vertex will form the minimum vertex cover. It is handled by Line 32 of Algorithm 3., vertex $v$ either (i) represents more than one vertex or (ii) is represented by more than one vertex or (iii) both. Hence, removing vertex $u$ and freezing $v$ is appropriate. Vertex $v$ is added to the vertex cover $S$. Importantly, due to the presence of recursive calls in Algorithm 4, the loop in Line 10 of Algorithm 3 is executed top-down. At this stage, a top down execution will facilitate removal or freezing of more vertices by the recursive calls of Algorithm 4 as compared to a bottom-up execution (Observation 1).
  • Lines 20 to 35 of Algorithm 3 freezes and removes the current “necessary” and “terminal” vertices, respectively: The execution of the loop in Line 20 will happen bottom-up. During $i^{th}$ iteration of the loop, if both the endpoints are neither frozen nor removed, then it implies that the endpoints represent each other only. Hence, we freeze the endpoint that is represented by more number of endpoints in $R$ and remove the other. Due to Observation 2, the decision to freeze the endpoint that is represented more in $R$ is valid. In case both the endpoints are represented by the same number of endpoints, tie is broken based on lexicographical ordering (as discussed at the beginning of Section 4). Note that represents table $R$ is continuously updated after each freeze or remove operation. Consequently, the represents list of each endpoint gets updated and it may become a necessary or a terminal vertex [^8] We stress on the word “current” due to the continuous updates to represents table $R$.. This is handled by Line 7 and 10 of Algorithm 4, respectively. Specifically, Line 7 of Algorithm 4 freezes an endpoint that represents a removed vertex. By induction, this is like removing terminal vertices and freezing its neighboring endpoints, but dependent on the current state of the represents table $R$. Similarly, Line 10 of Algorithm 4 removes an endpoint that does not represent any vertex based on the current state of the represents table $R$. This is a valid removal as the vertex had not been frozen yet and it does not represent any vertex in the latest iteration.
The above discussed sequential implementation of freezing “necessary” vertices, removing “terminal” vertices, and freezing and removing the current “necessary” and “terminal” vertices, respectively, ensures that the frozen vertices (equivalently, the vertices in vertex cover $S$) form the smallest vertex cover.
In summary, Algorithm 3 returns the smallest possible vertex cover derivable from the endpoints in $R$ given as input. Formally, variable $S$ in Line 9 of Algorithm 1 is the smallest vertex cover such that for all vertex covers $S^{\prime}\subseteq V^{\prime}$, $|S|\leq|S^{\prime}|$ where $V^{\prime}$ are the endpoints in $R$.
Lemma 8.
Given $m$ sets of endpoints $V^{\prime}\subseteq V$ of edges selected during maximal matching, there is at least one set $V^{\prime}$ that is a super-set of the set of vertices in the minimum vertex cover.
Proof.
Algorithm 3 (local minimization) finds the smallest vertex cover from a given set of vertices (Lemma 7). We now show that the input to Algorithm 3 consists of the following cases, each of which ensures that for every graph $G$, Algorithm 3 will find its minimum vertex cover:
  • perfect matching [^9] A perfect matching $M_{P}$ matches all the vertices of a graph. Hence, $|M_{P}|=\frac{m}{2}$. (all vertices): When Algorithm 2 (maximal matching) finds a perfect matching, all vertices will be added as endpoints in the represents table $R$. Hence, the smallest vertex cover that Algorithm 3 returns is indeed the minimum vertex cover.
  • maximum matching: There can be multiple maximum matching in a graph. Algorithm 1 uses one maximum matching. Hence, there are four possibilities:
    1. 1.
      endpoints of the maximum matching $E_{M}$ is a super set of the minimum vertex cover and Algorithm 2 traverses through $E_{M}$: This is a trivial case and Algorithm 3 returns the minimum vertex cover.
    2. 2.
      endpoints of the maximum matching $E_{M}$ is not a super set of the minimum vertex cover and Algorithm 2 traverses through $E_{M}$: This case implies that there is some other maximum matching $E^{\prime}_{M}$ whose endpoints are a super set of the minimum vertex cover. In such cases, Algorithm 2 may traverse through $E_{M}$ during an initial iteration. However, there will always be an iteration of BFS seeded on a vertex not an endpoint in $E_{M}$ that will eventually ensure that Algorithm 2 traverses through $E^{\prime}_{M}$, which implies that Algorithm 3 returns the minimum vertex cover.
    3. 3.
      endpoints of the maximum matching $E_{M}$ is a super set of the minimum vertex cover and Algorithm 2 does not traverse through $E_{M}$: This case may occur when the seed for BFS is not an endpoint of $E_{M}$. There are two sub cases here: (i) the graph consists of odd cycles and hence there is another maximum matching or maximal matching that Algorithm 2 traverses through and whose endpoints are a super set of minimum vertex cover. Here, Algorithm 3 returns the minimum vertex cover. (ii) the another maximum matching that Algorithm 2 traverses through is not a super set of minimum vertex cover and hence Algorithm 3 does not return the minimum vertex cover. In such a case, the loop in Algorithm 1 will continue to iterate over different seeds of BFS and there always exists one seed that is an endpoint of an edge in $E_{M}$. This implies that Algorithm 2 will eventually traverse through $E_{M}$ and Algorithm 3 will return the minimum vertex cover.
    4. 4.
      endpoints of the maximum matching $E_{M}$ is not a super set of the minimum vertex cover and Algorithm 2 does not traverse through $E_{M}$: This case, in principle, is equivalent to point 2 but the ordering of the seeds selected for BFS are reversed. During the initial iterations, Algorithm 2 not traversing through $E_{M}$ implies it traverses through another maximum matching or maximal matching whose endpoints is a super set of minimum vertex cover. In turn, this implies that Algorithm 3 returns the minimum vertex cover. For example, this happens when the input graph is a wheel graph. The maximum matching $E_{M}$ may only consist of edges on the boundary. However, there is another maximum matching $E^{\prime}_{M}$ that consists of an edge whose one endpoint is the center vertex of the wheel graph and hence, collectively, whose endpoints are a super set of the minimum vertex cover. Note that, by design, Algorithm 2 in wheel graphs will never traverse through $E_{M}$.
  • maximal matching: Depending on the seed of BFS selected in Algorithm 1, an iteration may result into Algorithm 2 selecting edges that form a maximal matching that is not necessarily a maximum matching (Lemma 3). This is why the third phase is called maximal matching. During such an iteration, there are two possibilities:
    1. 1.
      endpoints of the maximal matching in Algorithm 2 is a super set of the minimum vertex cover: This is a trivial case and Algorithm 3 returns the minimum vertex cover.
    2. 2.
      endpoints of the maximal matching in Algorithm 2 is a not super set of the minimum vertex cover: This case may occur even when the seed of BFS is a vertex that is an endpoint of an edge in maximum matching $E_{M}$. However, this seed is a terminal vertex (i.e. a vertex with degree = 1 in graph $G$). Hence, during one of the iterations of loop in Algorithm 1 when the the BFS is seeded on a non-terminal vertex that is an endpoint in $E_{M}$, one of the cases discussed in “maximum matching” will occur and Algorithm 3 will return the minimum vertex cover.
These cases complete the proof for this lemma. ∎
Lemma 9.
Algorithm 3 returns a minimum vertex cover.
Proof.
The proof follows due to a combination of Lemma 6, Lemma 7 and Lemma 8. Specifically, Lemma 6 proved that $S$ is a vertex cover, Lemma 7 proved that $S$ is the smallest vertex cover and Lemma 8 proved that there is at least one iteration where the input to Algorithm 3 consists of a super set of the minimum vertex cover and it returns a minimum vertex cover. Hence, as a combination of these lemmas, $S$ is indeed the minimum vertex cover. ∎
Overall, Lemma 9 means that if Algorithm 1 returns “Yes”, then the given instance of VC is a “Yes” instance. This completes the proof in the reverse direction. In turn, it completes the proof of Theorem 1.

### v2

# On Efficient Computation of DiRe Committees

@@ Abstract
The DiRe (Diverse + Representative) committee feasibility problem (a.k.a. the minimum vertex cover problem on unweighted undirected graphs) concerns the determination of the smallest size committee that satisfies the given constraints. Here, for this problem, we propose an algorithm that is an amalgamation of maximum matching, breadth-first search, maximal matching, and local minimization. We prove the algorithm terminates in polynomial-time. We conjecture the algorithm is an unconditional deterministic polynomial-time algorithm.
Preface: The DiRe committee feasibility problem (stated in the abstract) and the vertex cover problem on unweighted undirected graphs are equivalent (vertices == candidates; edges == candidate groups / voter populations’ approved candidates; for details, see Appendix A). Hence, for technical simplicity, we henceforth focus the discussion on the latter problem.
@@ 1 Introduction
Given an unweighted undirected graph (specifically, a 2-uniform hypergraph), the vertex cover of the graph is a set of vertices that includes at least one endpoint of every edge of the graph. Formally, given a graph $G=(V,E)$ consisting of a set of vertices $V$ and a collection $E$ of 2-element subsets of $V$ called edges, the vertex cover of the graph $G$ is a subset of vertices $S\subseteq V$ that includes at least one endpoint of every edge of the graph, i.e., for all $e\in E$, $e\cap S\neq\phi$. The corresponding computational problem of finding the minimum-size vertex cover (MVC) is NP-complete [^3] Strictly speaking, the decision version of the vertex cover problem is NP-complete whereas MVC itself (search version) is NP-hard. See Section 2.1 of [Kho19] for a lucid explanation delineating (a) search and decision problems and (b) NP-hardness and NP-completeness. [Coo71, Lev73, Kar72], which means that there is no known deterministic polynomial-time algorithm to solve MVC. Here, we conjecture an unconditional deterministic polynomial-time algorithm for MVC on unweighted simple connected graphs [^4] We subtly yet drastically switch the discussion from unweighted undirected graphs to unweighted simple connected graphs. For simplicity, we want to avoid having loops and/or unconnected components in the graph. In the context of this paper, this switch has no impact on the NP-completeness of the problem (Appendix C). Notwithstanding, in the case of the presence of loops, our algorithm will work (with minor modifications) if each loop is replaced by adding a dummy vertex and a corresponding edge. In the case of unconnected components, we can run the algorithm for each connected component independently and take a union of each of the minimum vertex covers to get the final minimum vertex cover..
@@ 3.3 Maximal Matching
Given a graph $G$, a matching $M$ is said to be maximal if for all other matching $^{\prime}, MM$$\not\subset M^{\prime}$.
@@
We may use Lemma 3 in the proof of correctness and explain why the third phase is called maximal matching and not maximum matching.
@@ 3.5 Summary
The algorithm we discovered is an amalgamation of the above-discussed phases. The sequential implementation of these phases ensures we get a minimum vertex cover cover. At a high-level, this is because: (i) Maximum matching and breadth-first search ensures that the edges selecting during the maximal matching phase follows a procedure as opposed to vanilla maximal matching where edges are selected randomly. (ii) Maximal matching implies we get a vertex cover. (iii) Local minimization ensures we get the smallest vertex cover. Overall, we conjecture that the combination of all these implies we get the minimum vertex cover.
Non-technical Comment: As discussed, the flow of the algorithm is as follows: maximum matching $\rightarrow$ breadth-first search $\rightarrow$ maximal matching $\rightarrow$ local minimization. However, the evolution of the algorithm happened in the following order: maximal matching $\rightarrow$ local minimization $\rightarrow$ breadth-first search $\rightarrow$ maximum matching. Indeed, eventually “prefixing” the algorithm with maximum matching helped us deal with the messy cycles, especially odd cycles. Recall that Blossom algorithm [Edm65] had to do “extra work” just to deal with odd cycles.
@@ 13 $S$ = $\mathtt{Local\_Minimization}$($R$)
Result: returns $R$ - Represents Table
2
3
4RR = a two-column table, Represents Table, that stores the endpoints of an edge selected during maximal matching and the corresponding vertices each endpoint represents
5 for each levellevel in $BFS_{level}$ do
@@ 20      If any vertex becomes edgeless in $G$, mark the vertex as visited in $BFS_{level}$
Result: returns $S$ - the smallest vertex cover
2
3
4SS = $\phi$
@@ 9    $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $v$, $\phi$)
14// The following for loop will traverse through the table $R$ top-down
15 for each rowrow in $R$ do
16 if if any one endpoint in rowrow is either frozen or removed then
@@
18 else if one endpoint $u$ in rowrow only represents another endpoint vertex $v$ in rowrow and $v$ represents more than one vertex then
@@ 20       $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $v$, $u$)
27// The following for loop will traverse through the table $R$ bottom-up
28 for each rowrow in $R$ do
@@ 32    $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $u$, $\phi$)
36      //at this point, both endpoints $u$ and $v$ in rowrow represent exactly one vertex, namely each other
@@ 44       $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $u$, $v$)
1
Data: Represents Table $R$, Vertex Cover $S$, vertex to be frozen freezefreeze, vertex to be removed removeremove
@@
2
3 Remove vertex removeremove and its represents list from $R$
4 Freeze vertex freezefreeze in $R$
5 Append vertex freezefreeze to $S$
6 Remove vertex freezefreeze from every represents list in $R$
7 Remove the represents list of vertex freezefreeze in $R$
8
9for each non-frozen and unremoved endpointendpoint in $R$ that represents removeremove do
10 $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, endpointendpoint, $\phi$)
@@
12
13for each non-frozen and unremoved endpointendpoint in $R$ that does not represent any vertex do
14 $R$, $S$ = $\mathtt{Freeze\_and\_Remove}$($R$, $S$, $\phi$, endpointendpoint)
@@
16
17return $R$, $S$
Algorithm 4 $\mathtt{Freeze\_and\_Remove}$($R$, $S$, freezefreeze, removeremove)
We conjecture the following:
Conjecture 1.